# =arxiv_version.qry

### 20 Dec 2023 v3

@@ 4.4.2 Model Representation Intervention
Apart from erasure, perturbation is another form of neuron-targeted intervention. One representative example is causal mediation analysis Vig et al. (2020), which measures the effect of a control variable on a response variable, mediated by an intermediate variable (or mediator). In machine learning, we can think of the input example as the control variable, the model output as the response variable, and an internal neuron as the mediator. Vig et al. (2020) use this framework to analyze gender bias in LMs. Through a case study on GPT-2, the authors show that gender bias is concentrated in a relatively small proportion of neurons, especially in the middle layers.
@@ 4.4.3 Strengths and Weaknesses
However, they also share a number of disadvantages. First, compared to other methods, counterfactual intervention is relatively more expensive in computational cost, normally requiring multiple forward passes and modifications to the model representation. Searching for the right targets to intervene in can also be costly. Second, as explained before, erasure-based intervention can result in nonsensical inputs or representations, which sometimes allow adversaries to manipulate the explanation Slack et al. (2020). Third, intervening in a single feature relies on the assumption that features are independent. Consider the sentence This movie is mediocre, maybe even bad Wallace, Gardner, and Singh (2020). If we mask out mediocre or bad individually, the predicted sentiment will probably not change much (still negative). Hence, an explanation method that relies on single-feature erasure might report that neither token is important for model prediction. However, such a method does not capture feature interactions, like the OR relationship between mediocre or bad here – as long as one of them is present, the sentiment is likely negative. More examples are provided in Shrikumar, Greenside, and Kundaje (2017). Addtionally, interventions are often overly specific to the particular example Wallace, Gardner, and Singh (2020). This calls for more insights into the scale of such explanations (i.e., if we discover a problem, is it only about one example or a bigger issue?) and general takeaways (i.e., what do we know about the model from this explanation?). Finally, counterfactual intervention may suffer from hindsight bias De Cao et al. (2020), which questions the foundation of counterfactual reasoning. Specifically, the fact that a feature can be dropped without influencing the prediction might not mean that the model “knows” that it can be dropped and has not used it in the original prediction. De Cao et al. (2020) illustrates this point with an intuitive example of the Reading Comprehension task, where a model is given a paragraph and a question, and should identify an answer span in the paragraph. Now, using counterfactual intervention, if we mask out everything except the answer in the paragraph, the model will for sure predict the gold span. Nonetheless, this does not imply that everything else is unimportant for the model’s original prediction. The issue again calls for a rethinking of the fundamental assumptions of counterfactual reasoning.
@@ 4.5.1 Explainable Architecture
While end-to-end NNs are a black-box, classic machine learning models like Decision Tree and linear regression have a highly interpretable reasoning mechanism. Drawing inspiration from them, researchers attempt to design neural models with more structural transparency while maintaining their performance.
@@
NMNs can be thought of as a variant of Neural-Symbolic Models (NSMs), a more generic concept loosely defined as neural models that integrate symbolic reasoning (Yi et al., 2018; Mao and Gan, 2019; Bogin et al., 2021, i.a.). Interested readers can refer to Hamilton et al. (2022) for a focused review on neuro-symbolic NLP.
Models with constraints are another example of explainable architectures. The idea is to incorporate constraints into neural networks from classic interpretable models, like generalized linear regression Alvarez Melis and Jaakkola (2018) and finite-state automata Schwartz, Thomson, and Smith (2018); Deutsch, Upadhyay, and Roth (2019); Jiang et al. (2020). Still, a major challenge lies in the trade-off between interpretability and performance.
@@ 4.5.2 Generating Explanations
Approaches in (i) suffer from a similar faithfulness challenge as predict-and-explain methods do, because the predictor can make its decision only based on the input using whatever reasoning mechanism, while totally ignoring the explanation. For example, Rajani et al. (2019) introduce a QA model that takes in a question, generates an explanation first, and then produces an answer based on both the question and the explanation. Another example is a variant of the NILE model Kumar and Talukdar (2020) previously discussed in explain-then-predict, which allows the predictor to look at both the premise-hypothesis pair and the explanation.
@@
In particular, we elaborate on a series of work on generating structured proofs for deductive reasoning. This is especially interesting since most previous work only aims at generating single-step explanations (e.g., a single sentence), but complex tasks would require a structured reasoning chain as explanations. As one of the earliest studies of this kind, Tafjord, Dalvi, and Clark (2021) develop ProofWriter, which takes in a set of premises and a hypothesis, and then decides if the hypothesis is true by providing a structured proof. Specifically, there are two versions of ProofWriter, all-at-once (generating the entire proof in one shot) and iterative (generating one step at a time). The all-at-once ProofWriter cannot guarantee that the proof is faithful, since it may not “believe” in the proof that it has generated. The iterative ProofWriter bridges this gap by deriving the proof one step at a time. All steps are already verified during generation. Therefore, it is faithful by construction. However, compared to the all-at-once version, it suffers from efficiency and input length limitations.
@@
With the advances in in-context learning Brown et al. (2020), recent work has started to explore the possibility of generating explanations with few-shot prompting. Earliest studies including Wiegreffe et al. (2022) and Marasović et al. (2022) find that Large Language Models (LLMs) like GPT-3 show potential for generating decently plausible free-text explanations with only a few examples, though the quality is still far from human-provided ones. On the other hand, Ye and Durrett (2022) find that such few-shot model-generated explanations are often non-factual (i.e., not correctly grounded in the input) and inconsistent (i.e., not entailing the prediction). Nonetheless, these explanations are still helpful for users to calibrate the confidence of model predictions, since explanations rated by humans as plausible are more likely to co-occur with accurate predictions.
@@
Another line of work under few-shot explanation generation is Chain-of-Thought-style (CoT) prompting, which is specifically effective for complex reasoning tasks like Math Word Problems and Multi-hop QA. Given a complex question $Q$, an LM is prompted to generate a reasoning chain $C$ along with the final answer $A$. Specifically, the prompt consists of a few instances of $(Q,C,A)$ triples as in-context exemplars. This allows pre-trained LLMs to solve unseen questions with much higher accuracy than standard prompting Brown et al. (2020), where the exemplars do not contain the reasoning chain $C$. We categorize existing CoT-style prompting methods into three types: all-at-once, ensemble-based, and modularized. All-at-once prompting means that the LM produces $C$ and $A$ as one continuous string, without any dependencies or constraints in between. Scratchpad Nye et al. (2021), CoT Wei et al. (2022), and “Let’s think step by step” Kojima et al. (2022), are all examples of this kind. Ensemble-based prompting is designed to overcome the locality issue of one-shot generation in previous methods by sampling multiple $(C,A)$ pairs and choosing the best answer via strategies like majority voting. Examples include Self-Consistent CoT Wang et al. (2022b), Minerva Lewkowycz et al. (2022), and DIVERSE Li et al. (2022), which differ mainly in the granularity of voting and the underlying LM. Modularized methods break down $Q$ into subproblems and then conquer them individually Hong et al. (2022); Creswell and Shanahan (2022); Qian et al. (2022); Jung et al. (2022). In particular, Least-to-Most Prompting Zhou et al. (2022) uses an LLM to first reduce the question to subquestions, and then sequentially answers them conditioned on its answers to previous subquestions. However, for all the above CoT-style prompting methods, the generated reasoning chain is entirely in NL, so there is no guarantee of faithfulness: the answer does not need to causally follow from the reasoning chain. To bridge this gap, recent work attempts to generate the reasoning chain in some Symbolic Language (SL) (e.g., Python) and calls an external solver (e.g., a Python interpreter) to derive the answer by deterministically executing the reasoning chain. In Program-of-Thought (PoT) Chen et al. (2022b) and Program-Aided Language Models (PAL) Gao et al. (2023), the reasoning chain is entirely in Python, which allows the same underlying LM to outperform vanilla CoT with NL by a large margin on a wide range of arithmetic and symbolic reasoning tasks. Lyu et al. (2023) propose Faithful CoT, which interleaves NL comments and SL programs for users to better understand and potentially interact with the model. These symbolic CoT prompting methods guarantee that the reasoning chain is a faithful explanation of how the model derives the answer. However, how the model generates the reasoning chain is still an opaque process, so there is no full interpretability of the entire pipeline.
@@ 4.5.3 Strengths and Weaknesses
In summary, self-explanatory models have several strengths. First, by definition, self-explanatory models provide built-in explanations, so there is no need for post-hoc explanations. Second, the form of explanation is flexible, e.g., model architecture, input features, natural language, or causal graphs. Third, it is possible to supervise the explainer with human-provided explanations. This is helpful for learning more plausible explanations, as well as encouraging the model to rely on desired human-like reasoning mechanisms instead of spurious cues. Finally, certain self-explanatory models (e.g., Tafjord, Dalvi, and Clark, 2021; Hong et al., 2022; Chen et al., 2022b; Gao et al., 2023; Lyu et al., 2023), are faithful by construction (we should be extra cautious about this claim, though).
@@ 5 Summary and Discussion
Backpropagation-based methods are faithfulness-driven by definition, but they still encounter theoretical challenges such as saturation, input sensitivity, and implementation variance. Subsequent variants have addressed these issues by considering the input alongside the gradient or by incorporating a baseline for comparison. Empirically, these methods are more frequently assessed for faithfulness, mainly through perturbation-based evaluation, than earlier methods. However, certain methods in this family are shown to be only doing partial input recovery, regardless of the model’s prediction. Meanwhile, much progress has been made on improving the plausibility of these methods, especially in reducing the visual noise in relevance maps.
@@ 6 Conclusion
In summary, similarity-based methods are not primarily driven by faithfulness and are rarely assessed on this basis. With respect to the analysis of model-internal structures, the early attempt to treat raw attention weights as explanations has faced strong criticism regarding faithfulness. However, there has been promising new progress in improving the faithfulness of attention by addressing these concerns. Backpropagation-based methods are intuitively faithfulness-motivated due to the nature of gradients, but theoretical and empirical evidence often points to various faithfulness issues, such as saturation and input sensitivity. Still, later variants in this family have shown gradual improvements in some of these aspects. Counterfactual intervention methods, rooted in causal inference, are also faithfulness-motivated. Nevertheless, certain types of intervention, such as perturbation, are more likely to be faithful than others like erasure, due to the OOD issue. There are also nuanced practical concerns such as feature independence and hindsight bias to consider. Self-explanatory models, which do not rely on any post-hoc explanation methods, often claim to be “faithful by construction”, yet many fall short due to obstacles like lack of label leakage. As such, we need to be extra cautious about such claims. In essence, when deciding which explanation methods to use in practice, we advocate for those that are intrinsically motivated and empirically validated in terms of faithfulness, while still remaining aware of the potential pitfalls highlighted above.

### 12 Jan 2024 v4

@@ 4.4.2 Model Representation Intervention
Apart from erasure, perturbation is another form of neuron-targeted intervention. One representative example is causal mediation analysis Vig et al. (2020), which measures the effect of a control variable on a response variable, mediated by an intermediate variable (or mediator). =In machine learning, we can think of the input example as the control variable, the model output as the response variable, and an internal neuron as the mediator. Vig et al. (2020) use this framework to analyze gender bias in LMs. Through a case study on GPT-2, the authors show that gender bias is concentrated in a relatively small proportion of neurons, especially in the middle layers.
@@ 4.4.3 Strengths and Weaknesses
However, they also share a number of disadvantages. First, compared to other methods, counterfactual intervention is relatively more expensive in computational cost, normally requiring multiple forward passes and modifications to the model representation. Searching for the right targets to intervene in can also be costly. Second, as explained before, erasure-based intervention can result in nonsensical inputs or representations, which sometimes allow adversaries to manipulate the explanation Slack et al. (2020). Third, intervening in a single feature relies on the assumption that features are independent. Consider the sentence This movie is mediocre, maybe even bad Wallace, Gardner, and Singh (2020). If we mask out mediocre or bad individually, the predicted sentiment will probably not change much (still negative). Hence, an explanation method that relies on single-feature erasure might report that neither token is important for model prediction. However, such a method does not capture feature interactions, like the OR relationship between mediocre or bad here – as long as one of them is present, the sentiment is likely negative. More examples are provided in Shrikumar, Greenside, and Kundaje (2017). Additionally, interventions are often overly specific to the particular example Wallace, Gardner, and Singh (2020). This calls for more insights into the scale of such explanations (i.e., if we discover a problem, is it only about one example or a bigger issue?) and general takeaways (i.e., what do we know about the model from this explanation?). Finally, counterfactual intervention may suffer from hindsight bias De Cao et al. (2020), which questions the foundation of counterfactual reasoning. Specifically, the fact that a feature can be dropped without influencing the prediction might not mean that the model “knows” that it can be dropped and has not used it in the original prediction. De Cao et al. (2020) illustrates this point with an intuitive example of the Reading Comprehension task, where a model is given a paragraph and a question, and should identify an answer span in the paragraph. Now, using counterfactual intervention, if we mask out everything except the answer in the paragraph, the model will for sure predict the gold span. Nonetheless, this does not imply that everything else is unimportant for the model’s original prediction. The issue again calls for a rethinking of the fundamental assumptions of counterfactual reasoning.
@@ 4.5.1 Explainable Architecture
While end-to-end NNs are a black-box, classic machine learning models like Decision Trees and linear regression have a highly interpretable reasoning mechanism. Drawing inspiration from them, researchers attempt to design neural models with more structural transparency while maintaining their performance.
@@
NMNs can be thought of as a variant of Neural-Symbolic Models (NSMs), a more generic concept loosely defined as neural models that integrate symbolic reasoning (Yi et al., 2018; Mao et al., 2019; Bogin et al., 2021, i.a.). Interested readers can refer to Hamilton_2022 for a focused review on neuro-symbolic NLP.
Models with constraints are another example of explainable architectures. The idea is to incorporate constraints into neural networks from classic interpretable models, like generalized linear regression Alvarez-Melis and Jaakkola (2018) and finite-state automata Schwartz, Thomson, and Smith (2018); Deutsch, Upadhyay, and Roth (2019); Jiang et al. (2020). Still, a major challenge lies in the trade-off between interpretability and performance.
@@ 4.5.2 Generating Explanations
Approaches in (i) suffer from a similar faithfulness challenge as predict-and-explain methods do, because the predictor can make its decision only based on the input using whatever reasoning mechanism, while entirely ignoring the explanation. For example, Rajani et al. (2019) introduce a QA model that takes in a question, generates an explanation first, and then produces an answer based on both the question and the explanation. Another example is a variant of the NILE model Kumar and Talukdar (2020) previously discussed in explain-then-predict, which allows the predictor to look at both the premise-hypothesis pair and the explanation.
@@
In particular, we elaborate on a series of studies on generating structured proofs for deductive reasoning. This is especially interesting since most previous work only aims at generating single-step explanations (e.g., a single sentence), but complex tasks would require a structured reasoning chain as explanations. As one of the earliest studies of this kind, Tafjord, Dalvi, and Clark (2021) develop ProofWriter, which takes in a set of premises and a hypothesis, and then decides if the hypothesis is true by providing a structured proof. Specifically, there are two versions of ProofWriter, all-at-once (generating the entire proof in one shot) and iterative (generating one step at a time). The all-at-once ProofWriter cannot guarantee that the proof is faithful, since it may not “believe” in the proof that it has generated. The iterative ProofWriter bridges this gap by deriving the proof one step at a time. All steps are already verified during generation. Therefore, it is faithful by construction. However, compared to the all-at-once version, it suffers from efficiency and input length limitations.
@@
With the advances in in-context learning Brown et al. (2020), recent work has started to explore the possibility of generating explanations with few-shot prompting. Earliest studies including Wiegreffe et al. (2022) and Marasovic et al. (2022) find that Large Language Models (LLMs) like GPT-3 show potential for generating decently plausible free-text explanations with only a few examples, though the quality is still far from human-provided ones. On the other hand, Ye and Durrett (2022) find that such few-shot model-generated explanations are often non-factual (i.e., not correctly grounded in the input) and inconsistent (i.e., not entailing the prediction). Nonetheless, these explanations are still helpful for users to calibrate the confidence of model predictions, since explanations rated by humans as plausible are more likely to co-occur with accurate predictions.
@@
Another line of work under few-shot explanation generation is Chain-of-Thought-style (CoT) prompting, which is specifically effective for complex reasoning tasks like Math Word Problems and Multi-hop QA. Given a complex question $Q$, an LM is prompted to generate a reasoning chain $C$ along with the final answer $A$. Specifically, the prompt consists of a few instances of $(Q,C,A)$ triples as in-context exemplars. This allows pre-trained LLMs to solve unseen questions with much higher accuracy than standard prompting Brown et al. (2020), where the exemplars do not contain the reasoning chain $C$. We categorize existing CoT-style prompting methods into three types: all-at-once, ensemble-based, and modularized. All-at-once prompting means that the LM produces $C$ and $A$ as one continuous string, without any dependencies or constraints in between. Scratchpad Nye et al. (2021), CoT Wei et al. (2022), and “Let’s think step by step” Kojima et al. (2022), are all examples of this kind. Ensemble-based prompting is designed to overcome the locality issue of one-shot generation in previous methods by sampling multiple $(C,A)$ pairs and choosing the best answer via strategies like majority voting. Examples include Self-Consistent CoT Wang et al. (2022b), Minerva Lewkowycz et al. (2022), and DIVERSE Li et al. (2022), which differ mainly in the granularity of voting and the underlying LM. Modularized methods break down $Q$ into subproblems and then conquer them individually Hong et al. (2022); Creswell and Shanahan (2022); Qian et al. (2022); Jung et al. (2022). In particular, Least-to-Most Prompting Zhou et al. (2022a) uses an LLM to first reduce the question to subquestions, and then sequentially answers them conditioned on its answers to previous subquestions. However, for all the above CoT-style prompting methods, the generated reasoning chain is entirely in NL, so there is no guarantee of faithfulness: the answer does not need to causally follow from the reasoning chain. To bridge this gap, recent work attempts to generate the reasoning chain in some Symbolic Language (SL) (e.g., Python) and calls an external solver (e.g., a Python interpreter) to derive the answer by deterministically executing the reasoning chain. In Program-of-Thought (PoT) Chen et al. (2022b) and Program-Aided Language Models (PAL) Gao et al. (2022), the reasoning chain is entirely in Python, which allows the same underlying LM to outperform vanilla CoT with NL by a large margin on a wide range of arithmetic and symbolic reasoning tasks. Lyu et al. (2023) propose Faithful CoT, which interleaves NL comments and SL programs for users to better understand and potentially interact with the model. These symbolic CoT prompting methods guarantee that the reasoning chain is a faithful explanation of how the model derives the answer. However, how the model generates the reasoning chain is still an opaque process, so there is no full interpretability of the entire pipeline.
@@ 4.5.3 Strengths and Weaknesses
In summary, self-explanatory models have several strengths. First, by definition, self-explanatory models provide built-in explanations, so there is no need for post-hoc explanations. Second, the form of explanation is flexible, e.g., model architecture, input features, natural language, or causal graphs. Third, it is possible to supervise the explainer with human-provided explanations. This is helpful for learning more plausible explanations, as well as encouraging the model to rely on desired human-like reasoning mechanisms instead of spurious cues. Finally, certain self-explanatory models (e.g., Tafjord, Dalvi, and Clark, 2021; Hong et al., 2022; Chen et al., 2022b; Gao et al., 2022; Lyu et al., 2023), are faithful by construction (we should be extra cautious about this claim, though).
@@ 5 Summary and Discussion
Backpropagation-based methods are faithfulness-driven by definition, but they still encounter theoretical challenges such as saturation, input sensitivity, and implementation variance. Subsequent variants have addressed these issues by considering the input alongside the gradient or by incorporating a baseline for comparison. Empirically, these methods are more frequently assessed for faithfulness, mainly through perturbation-based evaluation, than earlier methods. However, certain methods in this family are shown to be only doing partial input recovery, regardless of the model’s prediction. Meanwhile, much progress has been made in improving the plausibility of these methods, especially in reducing the visual noise in relevance maps.
@@ 6 Conclusion
In summary, similarity-based methods are not primarily driven by faithfulness and are rarely assessed on this basis. For the analysis of model-internal structures, the early attempt to treat raw attention weights as explanations has faced strong criticism regarding faithfulness. However, there has been promising new progress in improving the faithfulness of attention by addressing these concerns. Backpropagation-based methods are intuitively faithfulness-motivated due to the nature of gradients, but theoretical and empirical evidence often points to various faithfulness issues, such as saturation and input sensitivity. Still, later variants in this family have shown gradual improvements in some of these aspects. Counterfactual intervention methods, rooted in causal inference, are also faithfulness-motivated. Nevertheless, certain types of intervention, such as perturbation, are more likely to be faithful than others like erasure, due to the OOD issue. There are also nuanced practical concerns such as feature independence and hindsight bias to consider. Self-explanatory models, which do not rely on any post-hoc explanation methods, often claim to be “faithful by construction”, yet many fall short due to obstacles like lack of label leakage. As such, we need to be extra cautious about such claims. In essence, when deciding which explanation methods to use in practice, we advocate for those that are intrinsically motivated and empirically validated in terms of faithfulness, while still remaining aware of the potential pitfalls highlighted above.
@@
Acknowledgements.
This research is based upon work supported in part by the Air Force Research Laboratory (contract FA8750-23-C-0507), the DARPA KAIROS Program (contract FA8750-19-2-1004), the IARPA HIATUS Program (contract 2022-22072200005), and the NSF (Award 1928631). Approved for Public Release, Distribution Unlimited. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of AFRL, DARPA, IARPA, NSF, or the U.S. Government.
[…]